OpenStudy (anonymous):

Medal and Fan if you help me on this!! 34x(3.54+8.19x-12.87)+87=9.68x-24(3+8.92)+3x

OpenStudy (mtalhahassan2):

hey

OpenStudy (anonymous):

hey! umm do you think you can help me on this? im having trouble...

OpenStudy (mtalhahassan2):

yeah I can

OpenStudy (mtalhahassan2):

278.46 x square - 317.22 x = -360.4

OpenStudy (mtalhahassan2):

This is the answer which I get

OpenStudy (mtalhahassan2):

Is it correct?

OpenStudy (georgesyip):

... a bit weird but this is all i could get to... 34x(3.54+8.19-12.87)+87=9.68X+24(3+8.92)+3x 34x(3.54+8.19x-12.87)=9.68x+286.08+3x+87 120.36x+278.36x^2-437.58x=12.68x+286.08+87 120.36x+278.46x^2-437.58x=12.68x=373.08 278.46x^2-317.22=12.68x+12.68x+373.08

OpenStudy (anonymous):

thanks a lot guys i wish i could give u all medals :D

OpenStudy (georgesyip):

278.46x^2-12.68x=690.3

OpenStudy (georgesyip):

yeah sorry if this doesnt help you much

OpenStudy (anonymous):

its ok! it helps thanks so much~

OpenStudy (mtalhahassan2):

You welcome

OpenStudy (anonymous):

thx to both of u guys :D